OpenStudy (anonymous):

Which ordered pairs in the form (x, y) are solutions to the equation 3x – 4y = 21? Choose all answers that are correct. A. (−3, 3) B. (−1, −6) C. (7, 0) D. (11, 3)

OpenStudy (anonymous):

Thanks @pooja195

pooja195 (pooja195):

Take your options and plug them into the given equation 3x-4y=21 A. \[\huge~\rm~3(-3)-4(3)=21 \] Is this true or false?

OpenStudy (anonymous):

hold on one sec let me work out the problem

OpenStudy (anonymous):

yes it's true right

pooja195 (pooja195):

-3 *3=?

OpenStudy (anonymous):

ohhhh false it's negative right

pooja195 (pooja195):

Correct :) ok so B. \[\huge~\rm~3(-1)-4(-6)=21\] true or false

OpenStudy (anonymous):

Yeah this one is true

OpenStudy (anonymous):

and thanks for the medal!!!

pooja195 (pooja195):

Well done :-) C \[\huge~\rm~3(7)-4(0)=21\] true or false

OpenStudy (anonymous):

True

pooja195 (pooja195):

Good now the last one :-) Well done :-) D \[\huge~\rm~3(11)-4(3)=21\] true or false

OpenStudy (anonymous):

FALSE sike im just kidding TRUE

pooja195 (pooja195):

xD Alright that should answer your question ^_^

OpenStudy (anonymous):

THANKS SO MUCH YOUR SO SMART AND ALOT OF HELP IM GLAD U DIDN'T GIVE ME THE DIRECT ANSWER YOUR SOOO SMART I WISH I COULD GIVE YOU 80 MEDALS!!!!!!

jaynator495 (jaynator495):

GAH

jaynator495 (jaynator495):

I was to slow again xD

OpenStudy (anonymous):

DONT WORRY I MIGHT HAVE MORE

pooja195 (pooja195):

haha yw I love to see people learn ^_^

jaynator495 (jaynator495):

Pooja im helping again xD

jaynator495 (jaynator495):

got bored stairing at os feedback so ._>

OpenStudy (anonymous):

lol

pooja195 (pooja195):

Good :-)